A town recently dismissed 7 employees in order to meet their new budget reductions. The town had 4 employees over 50 years of age and 16 under 50. If the dismissed employees were selected at random, what is the probability that no more than 1 employee was over 50? Express your answer as a fraction or a decimal number rounded to four decimal places.

Please show how to do in Excel!
Answer #1

Total employees = 4 + 16 = 20

P(No more than 1 employee was over 50)

= P(0 from over 50; 7 from under 50) + P(1 from over 50; 6 from under 50)

= combin(16, 7) / combin(20, 7) + (combin(4, 1) * combin(16, 6)) / combin(20, 7) [Excel Formula]

= 0.5608